OpenStudy (anonymous):

What is the value of the expression |x| + |y + z| when x = –6, y = –3, and z = –5?

OpenStudy (anonymous):

ill give medal

OpenStudy (anonymous):

|-6|=6 |-3-5|=|-8| |-8|=8 So... 6+8=14

OpenStudy (anonymous):

|x| + | y + z| when x = -6, y = -3, and z = -5 keep in mind, a;; absolute values are POSITIVE |-6| + |-3 + (-5) | = 6 + | -3-5| = 6 + | -8 | = 6 + 8 = 14 Lorena is correct

OpenStudy (anonymous):

typo...absolute values are positive

OpenStudy (anonymous):

what do i typo any hot girl help me